OpenStudy (erinweeks):

Use mathematical induction to prove that the statement is true for every positive integer n.

OpenStudy (erinweeks):

2 is a factor of n2 - n + 2

OpenStudy (perl):

basis case n=1 2 is a factor of 1^2 - 1 + 2 ? yes because 1^2 -1 + 2 = 2 if 2 is a factor of k^2 -k + 1 is it true that 2 is a factor of (k+1)^2 -(k+1) + 2 ? thats what we have to prove

OpenStudy (perl):

ok?

OpenStudy (erinweeks):

okay sorry my computer shut down due to updates! @perl

OpenStudy (perl):

hi

OpenStudy (erinweeks):

okay so how do we finish this?

OpenStudy (perl):

so we need to prove the following statement: "if 2 is a factor of k^2 -k + 1 for some integer k *then* is it true that 2 is a factor of (k+1)^2 -(k+1) + 2 "

OpenStudy (perl):

lets look at (k+1)^2 - (k+1) + 2 , expand that

OpenStudy (usukidoll):

proofs are hard :(

OpenStudy (erinweeks):

now let me ask quick why did you expand it?

OpenStudy (erinweeks):

@perl ?

OpenStudy (usukidoll):

I think he's away...

OpenStudy (perl):

we want to use the fact that k^2 - k + 2 is divisible by 2

OpenStudy (usukidoll):

oh nevermind...

OpenStudy (perl):

at some point

OpenStudy (erinweeks):

OK

OpenStudy (erinweeks):

okay than what

OpenStudy (perl):

(k+1)^2 - (k+1) + 2 = k^2 + 2k + 1 - (k+1) + 2 = k^2 + 2k + 1 - k - 1 + 2 = k^2 +2k -k + 1 -1 + 2 =k^2 + 2k - k + 2 = k^2 -k + 2 + 2k = (k^2 - k + 2) + 2k, but we know that (k^2 -k + 2) is divisible by 2 since we assumed that

OpenStudy (perl):

so we have two numbers both of which are divisible by 2. therefore the sum will be divisible by 2. and we are done

OpenStudy (erinweeks):

you kinda lost me now.

OpenStudy (perl):

we assumed that k^2 + -k + 2 is divisible by 2. using this assumption we can prove that (k+1)^2 - (k+1) + 2 is divisible by 2

OpenStudy (erinweeks):

okay so basically (k+1)^2 -(k+1) +2 would be my answer i dont have to put a # in for k

OpenStudy (perl):

you know the statement is true for n=1. and we prove that if the statement is true for n=k, then the statement is true for n=k+1. so the statment is true for n=1, then since it is true for n=1 it is true for n= 1+1 = 2. now since it is true for n=2 it is true for n= 2 + 1 = 3, etc

OpenStudy (erinweeks):

well am i suppose to just write that one equation though

OpenStudy (erinweeks):

Im confused on what to put/

OpenStudy (erinweeks):

????

OpenStudy (usukidoll):

he's working on more than one problem on os apparently

OpenStudy (erinweeks):

@UsukiDoll do you know exactally what i have to put

OpenStudy (usukidoll):

>_________<!

OpenStudy (usukidoll):

I haven't taken a proof writing class, so no sorry...but I would put what perl mentioned

OpenStudy (erinweeks):

i know but would it be the equation i wrote..

OpenStudy (usukidoll):

perhaps... do you know spherical coordinates?